Find P.D. from B.D. using Stirling’s approximation.



1
Expert's answer
2022-04-17T23:23:51-0400

According to the Stirling's approximation:

"n!\approx\sqrt{2\pi n}(\frac{n}{e})^n" for large n


The mean of the binomial distribution "\mu=np\Rarr p=\frac{\mu}{n}"


"P(x=k)=C_n^kp^k(1-p)^{n-k}=\\
=\frac{n!}{k!(n-k)!}(\frac{\mu}{n})^k(1-\frac{\mu}{n})^{n-k}\approx\\
\approx\frac{\sqrt{2\pi n}(n/e)^n}{\sqrt{2\pi(n-k)}((n-k)/e)^{n-k}k!}\frac{\mu^k}{n^k}\frac{(1-\mu/n)^n}{(1-\mu/n)^{k}}"


Let's use some approximations for large n:

"(1-\frac{\mu}{n})^n\approx e^{-\mu}\\
(1-\frac{\mu}{n})^k\approx1\\
\sqrt{\frac{n}{n-k}}\approx1\\
\frac{(n/e)^n}{((n-k)/e)^{n-k}}\approx\frac{e^kn^k(1-k/n)^k}{(1-k/n)^n}\approx\frac{e^{-k}n^k}{e^{-k}}=n^k"


"P(k)\approx n^k\frac{\mu^k}{n^kk!}e^{-\mu}=\frac{\mu^ke^{-\mu}}{k!}" - Poisson's distribution.